Construct a 95% confidence interval for the population standard deviation o of a random sample of 15 men who have a mean weight of 165.2 pounds with a standard deviation of 13.5 pounds. Assume the population is normally distributed.
Answer #1

Solution :

Given that,

Sample size = n = 15

Z/2 = 1.96

Margin of error = E = Z/2* ( /n)

= 1.96 * (13.5 / 15)

Margin of error = E = 6.8

At 95% confidence interval estimate of the population mean is,

- E < < + E

165.2 - 6.8 < < 165.2 + 6.8

158.4 < < 172.0

(158.4 , 172.0)